You are considering a new product launch. The project will cost $750,000, have a four-year life, and have no salvage value; depreciation is straight-line to zero. Sales are projected at 180 units per year; price per unit will be $18,500, variable cost per unit will be $11,400, and fixed costs will be $522,000 per year. The required return on the project is 12 percent, and the relevant tax rate is 28 percent.

a) Based on your experience, you think the unit sales, variable cost, and fixed cost projections given here are probably accurate to within ±13 percent. What are the best and worst cases for these projections? What is the base-case NPV? What are the best-case and worst-case scenarios?

b) Evaluate the sensitivity of your base-case NPV to changes in fixed costs.

Please show all steps for credit.
Answer #1

Base Case NPV:

A B H. M | 1 Year Operating profit (Sales unit*(price-VC)-FC Depreciation Taxable income Tax (1*28%) PAT (1-) 0 59 omto 00 In

For Best Case NPV : Sales=180*1.13=203 (Approx), Price=18500*1.13=$20905 and VC= $11400*(1-.13)=9918, FC= 522000*(1-.13)=$454140

Best Case NPV Calculation: Year Operating profit (Sales unit* (price-VC)-FC Depreciation Taxable income Tax (1*28%) PAT (1-1)

For Worst case NPV:

Sales=180*(1-.13)=157 (Approx), Price=18500*(1-.13)=$16095 and VC= $11400*(1+.13)=12882, FC= 522000*(1+.13)=$589860

Worst Case NPV Calculation: Year 0 Initial Cost Sales unit per year Price per unit VC per unit Fixed Cost 750,000 157 16,095

b) If you increase FC of base case by 1%, the NPV= $1051335, calculation is given below:

C E B Base Case NPV Calculation: Operating profit (Sales unit*(price-VC)-FC 2 Year Sales unit per year Price per unit VC per

Hence, Sensitivity of NPV=(1062751-1051335)/1062751= 1.074%